OpenStudy (anonymous):

logx8+logx4=5 : x is base solve for x

OpenStudy (anonymous):

\[ \log_x8+\log_x5=\log_x(8\times 5) \]

OpenStudy (anonymous):

Which means \(x^5=8\times 5\)

OpenStudy (campbell_st):

Slight typo should read x^5 = 8*4

OpenStudy (anonymous):

Oh yeah.

OpenStudy (anonymous):

the answer is 2